A truck travels some distance with the constant speed of 30kmh -1 and return back with the constant speed of 40kmh-1. Calculate the average speed of truck.

Grade:9

2 Answers

Arun
5 years ago
Dear Ashish
 
Avg speed = distance /time
 
 = x + x / (x/30 + x/40) = 2x /70x/1200
 
 = 2400x /70 x = 34.2855 km/hour
Kushal Chaudhari
5 years ago
The average speed is =total distance /total time 
Distance from starting to end =x 
Then total distance =2x
Since Speed=distance / time 
Therefore Time =distance /speed
Time from start to end point = x/30
Time from end to start point =x40
Total time taken =(x/30)+(x/40)
Average speed = total distance /total time
=2x/(x/30)+(x/40)
=2x/7x/120
=2x *120/7x
=240x/7x
=34.2857143 km /hour